OpenStudy (jaxon):

how do I find the slope of - x -y = -3

OpenStudy (amistre64):

most basic way to determine slope is see how far the y moves when you move x by 1

OpenStudy (aroub):

the slope is already written y=mx+b , -y=x-3 , y=-x+3 so the slope is -1

OpenStudy (amistre64):

-(0) -y = -3 ; when x=0; y=3 -(1) -y = -3; when x=1; y=2 y goes from 3 to 2; thats a direction of -1

OpenStudy (anonymous):

first keep it simple..like x+y-3=0 so slope=-a/b hence slope=-1

OpenStudy (jaxon):

i just cant seem to grasp this!

OpenStudy (amistre64):

there are a few ways to come up with the slope; but the most basic is to simply find the distance y moves when x moves by 1 ....

OpenStudy (jaxon):

but what do you mean? when Y moves? Sorry I am really struggling

OpenStudy (amistre64):

slope is how we measure how far something leans away from vertical

OpenStudy (amistre64):

when we graph an equation, we define slope as how far a point moves away from(or towards) the y axis

OpenStudy (amistre64):

it is the change in your y values with respect to a single change in your x value.